I am trying to solve the problem h(x) = x+17 when h equals h(y+9). So far I have solved for the equation using h(0), h(-1), H(-11), etc. but am struggling with how to solve when a y is involved. Do I solve for y first?
Answer by Theo(13342)

your equation is h(x) = x+17
you want to solve for h(y+9).
this is no different than when you solved for h(-11).
when you did that, you replaced x with -11 to get h(-11) = (-11) + 17 which gave you an answer of h(-11) = 6.
when you want to find h(y+9), then you replace x with (y+9) to get:
h(y+9) = (y+9)+17 = y+26
h(y+9) = 6+26
that's your answer.
you do exactly the same thing as when you solved for h(0) or h(-1) or h(-11).
you replace x with whatever you want to get the h of.
in this case h(y+9) = (y+9) + 17 = y + 26